# Efficient Exact Algorithms for Minimum Covering of Orthogonal Polygons with Squares

### v1

@@ 6.2 Fixing the Construction
The issue in the NP-hardness reduction in [3] arose as the clauses require more number of squares when the literals of a clause do not all evaluate to true even if this assignment sets the clause to true.
@@
The construction due to [3] starts by defining even-lines (odd-lines) as horizontal lines with an even (odd) y-coordinate or a vertical line with an even (odd) x-coordinate. The construction of variable gadgets and wires follow the following properties.
@@
  • The wires connecting a variable gadget for a variable $x$, to a junction for a clause $c$, where $x$ appears non-negated in $c$ — connect to the variable gadget (as well as the junction) horizontally, and along two consecutive odd lines (i.e. an even line passes through the middle of such a connection). Refer to diagrams in  [3].
  • The wires connecting a variable gadget for a variable $x$, to a junction for a clause $c$, where $x$ appears negated in $c$ — connect to the variable gadget (as well as the junction) horizontally, and along two consecutive even lines (i.e. an odd line passes through the middle of such a connection). Refer to diagrams in  [3].
  • If a variable gadget for a variable $x$ is covered with a true-covering, then all wires connecting it to a junction for a clause $c$ where $x$ appears non-negated, will be covered in a way such that half a square protrudes out of destination and hence covers a part of the junction. Otherwise, for a false-covering, the entire junction would remain uncovered, even when the variable gadget and the wire are covered. Refer to diagrams in  [3].
  • If a variable gadget for a variable $x$ is covered with a false-covering, then all wires connecting it to a junction for a clause $c$ where $x$ appears negated, will be covered in a way such that half a square protrudes out of destination and hence covers a part of the junction. Otherwise, for a true-covering, the entire junction would remain uncovered, even when the variable gadget and the wire are covered. Refer to diagrams in  [3].
We construct new junction gadgets along with these constructions of wires and variable gadgets in [3].
@@
We construct new junction gadgets as shown in Figure 15. Wires connect to these junction gadgets in the same way as described in [3].
@@
We now investigate minimal coverings of squares for each of these gadgets.
@@
Now, we complete the reduction from Planar 3-CNF (refer to [8]).
@@
We consider an instance (a CNF boolean formula) $\psi$ of Planar 3-CNF. We construct variable gadgets (as in [3]) for each variable in $\psi$, and set up junctions ( Figure 15) for each clause in $\psi$ (we are working with $\psi$ directly, instead of working with $\phi=\neg\psi$ as done in [3]) and connect them with wires (as in [3]). Let $V+W$ be the squares required to cover all variable gadgets and wires.
Now, we analyse the three cases like before. Let $j$ be the number of junctions.
  • Case I, $\psi$ is true for all assignments (satisfiable). Since $\psi$ is in CNF, all assignments must render all or-clauses as true (i.e. at least one literal in each clause is true). Therefore from Lemma 6.1, for all minimum covering of the variable gadgets, coverings all junctions would require a minimum of only $28$ valid squares. This means the minimum number of squares needed to cover such an instance is equal to $V+W+28j$.
  • Case II, $\psi$ is false for all assignments (not satisfiable). Since $\psi$ is in CNF, all assignments must render at least one or-clause as false (i.e. at all literals in at least one clause is false). Therefore from Lemma 6.1, at least one junction requires a minimum of $29$ valid squares, where as others require at least $28$ valid squares for being covered. This means the minimum number of squares needed to cover such an instance is strictly more than $V+W+28j$.
  • Case III, $\psi$ is true for some assignments and false for some (satisfiable). Consider any such assignment such that $\psi$ is true and cover the variable clauses accordingly. Since $\psi$ is in CNF, this assignments must render all or-clauses as true (i.e. at least one literal in each clause is true). Therefore from Lemma 6.1, for such a minimum covering of the variable gadgets, coverings all junctions would require $28$ squares. This means the minimum number of squares needed to cover such an instance is equal to $V+W+28j$.
This implies that $\psi$ is satisfiable if and only if the reduced instance has a minimum covering with exactly $V+W+28j$ squares. On the other hand, if $\psi$ is not satisfiable, then the minimum number of squares required to cover the reduced instance is strictly more than $V+Q+28j$. Since Planar 3-CNF is NP-hard, OPCSH must also be NP-hard. Finally, as any certificate can be verified in polynomial time using Algorithm 1, OPCSH is NP-complete.
@@
The problem of covering orthogonal polygons with holes using minimum number of squares (OPCSH) is NP-complete (when the input is the set of all $N$ lattice points inside the orthogonal polygon).
Moreover, as both the answer $V+W+28j$ and the numerical value of the coordinates of each of the points is linear (polynomial) in the number $N$ of points inside the polygon, the problem is strongly NP-complete. Hence we can conclude the following.
@@
  • The problem of finding the minimum square covering of polygons, where the squares can take side lengths from a set $\Lambda\subset\mathbb{N}$ is NP-complete, even if $|\Lambda|=2$.
Proof 6.6.
The results follow from the observation that the reduced instance from Planar 3-CNF only contains maximal squares of size $2\times 2$ and $4\times 4$, and from Observation 3.
@@
Proof 6.8.
Theorem 6.3 immediately gives us that OPCSH is NP-hard when the input is the $n$ vertices.
@@
  • If all rec-packs lie in the polygon, we can use Algorithm 1 to check if they form a complete covering of the polygon. Algorithm 1 does generalize to polygons with holes as well.
@@ 7 Conclusion
In this paper, we answer an open problem in [3], of whether OPCS has an exact algorithm running in time polynomial in the number $n$ of vertices of the input orthogonal polygon without holes. Our exact algorithm for OPCS runs in $\mathcal{O}(n^{14})$ time. We further optimize the running time for orthogonal polygons with $n$ vertices and a small number of knobs $k$, by designing a recursive algorithm with running time $\mathcal{O}(n^{2}+k^{14}\cdot n)$. This gives us an $\mathcal{O}(n^{2})$ algorithm for solving OPCS on orthogonally convex polygons. We also correct some errors in the NP-hardness reduction for OPCSH given in [3]. A natural future direction is to study the problem with respect to covering of orthogonal polygons with other geometric objects like triangles, line segments, orthogonally convex polygons etc. and try to obtain an algorithm whose running time is polynomial in the number of vertices of the input polygon and not on the total number of interior points of the input polygon. Another interesting question directly related to our current work would be to find tight bounds for the number of rec-packs required to produce a minimum covering of an orthogonal polygon; our paper only proves an upper bound of $\mathcal{O}(n^{2})$ rec-packs required to produce a minimum covering of an entire polygon.

### v2

@@ 7.2 Fixing the construction
The issue in this claimed NP-hardness reduction in arose as the clauses require more number of squares when the literals of a clause do not all evaluate to true even if this assignment sets the clause to true.
@@
The construction due to Aupperle et. al [4] starts by defining even-lines (odd-lines) as horizontal lines with an even (odd) y-coordinate or a vertical line with an even (odd) x-coordinate. The construction of variable gadgets and wires follow these properties:
@@
  • The wires connecting a variable gadget for a variable $x$, to a junction for a clause $c$, where $x$ appears non-negated in $c$ — connect to the variable gadget (as well as the junction) horizontally along two consecutive odd lines (i.e. an even line passes through the middle of such a connection). Please refer to diagrams of the original construction [4].
  • The wires connecting a variable gadget for a variable $x$, to a junction for a clause $c$, where $x$ appears negated in $c$ — connect to the variable gadget (as well as the junction) horizontally along two consecutive even lines (i.e. an odd line passes through the middle of such a connection). Please refer to diagrams of the original construction [4].
  • If a variable gadget for a variable $x$ is covered with a true-covering, then all wires connecting it to a junction for a clause $c$ where $x$ appears non-negated, will be covered in a way such that half a square protrudes out of destination and hence covers a part of the junction. Otherwise, for a false-covering, the entire junction would remain uncovered, even when the variable gadget and the wire are covered. Please refer to diagrams of the original construction [4].
  • If a variable gadget for a variable $x$ is covered with a false-covering, then all wires connecting it to a junction for a clause $c$ where $x$ appears negated, will be covered in a way such that half a square protrudes out of destination and hence covers a part of the junction. Otherwise, for a true-covering, the entire junction would remain uncovered, even when the variable gadget and the wire are covered. Please refer to diagrams of the original construction [4].
We will use these construction for variable gadgets as it is, but we provide novel constructions for junction gadgets as mentioned earlier.
@@
We construct new junction gadgets as shown in Figure 17. Wires connect to these junction gadgets in the same way as described in the original construction [4].
@@
We now investigate minimal coverings of squares for each of these gadgets. For the remaining paper, we denote by the term ‘literal’, a variable or its negation that appears in a clause.
@@
Now, we complete the reduction from Planar 3-CNF [15].
@@
We consider a CNF boolean formula $\psi$, which is an instance of Planar 3-CNF. We construct variable gadgets for each variable in $\psi$, and set up junctions (Figure 17) for each clause in $\psi$ and connect them with wires. Let $V+W$ be the squares required to cover all variable gadgets and wires. Note that we are working with $\psi$ directly, instead of working with $\phi=\neg\psi$ as done by Aupperle et. al [4].
Now, we analyse the three cases like before. Let $j$ be the number of junctions. Note that, due to Lemma 7.1, the minimum number of squares required to cover the reduced instance is at least $V+W+28j$.
  • Case I, $\psi$ is true for all assignments (satisfiable). Since $\psi$ is in CNF, all assignments must render all or-clauses as true; i.e. at least one literal in each clause is true. Therefore from Lemma 7.1, for all minimum covering of the variable gadgets, coverings all junctions would require a minimum of only $28$ valid squares. This means the minimum number of squares needed to cover such an instance is equal to $V+W+28j$.
  • Case II, $\psi$ is false for all assignments (not satisfiable). Since $\psi$ is in CNF, all assignments must render at least one or-clause as false; i.e. all literals in at least one clause is false. Therefore from Lemma 7.1, at least one junction requires a minimum of $29$ valid squares, whereas others require at least $28$ valid squares for being covered. This means the minimum number of squares needed to cover such an instance is strictly more than $V+W+28j$.
  • Case III, $\psi$ is true for some assignments and false for some (satisfiable). Consider any assignment such that $\psi$ is true. Since $\psi$ is in CNF, this assignments must render all or-clauses as true; i.e. at least one literal in each clause is true. Consider the square covering corresponding to this assignment. Due to Lemma 7.1, for such a minimum covering of the variable gadgets, coverings all junctions would require exactly $28$ squares. This means the minimum number of squares needed to cover such an instance is equal to $V+W+28j$.
This implies that $\psi$ is satisfiable if and only if the reduced instance has a minimum covering with exactly $V+W+28j$ squares. On the other hand, if $\psi$ is not satisfiable, then the minimum number of squares required to cover the reduced instance is strictly more than $V+Q+28j$. Since Planar 3-CNF is NP-hard, OPCSH must also be NP-hard. Finally, as any certificate can be verified in polynomial time (Lemma 5.1), OPCSH is NP-complete.
@@
The problem of covering orthogonal polygons with holes using minimum number of squares (OPCSH) is NP-complete, where the input is the set of all $N$ lattice points inside the orthogonal polygon.
Moreover, as both the answer $V+W+28j$ and the numerical value of the coordinates of each vertex is linear in the number $N$ of lattice points inside the polygon, the problem is strongly NP-complete, and hence cannot have a fully polynomial time approximation scheme (FPTAS) [11, 18].
@@
  • The problem of finding the minimum square covering of polygons, where the squares can take side lengths from a set $\Lambda\subseteq\mathbb{N}$ is NP-complete, even if $|\Lambda|=2$.
Proof 7.6.
The results follow from the observation that the reduced instance from Planar 3-CNF only contains maximal squares of size $2\times 2$ and $4\times 4$, and from Observation 2.
@@
Proof 7.8.
Theorem 7.3 immediately gives us that OPCSH is NP-hard when the input is the $n$ vertices, as the reduced instance of orthogonal polygon has $N=\Theta(n)$ lattice points inside.
@@
  • If all rec-packs lie in the polygon, we can use Lemma 5.1 to check if they form a complete covering of the polygon. Note that Lemma 5.1 generalizes to polygons with holes as well.
@@ 8 Conclusion
In this paper, we answer the open problem, initially posed by Aupperle et. al [4], asking whether OPCS has an exact algorithm running in time polynomial in the number $n$ of vertices of the input orthogonal polygon without holes. Our exact algorithm for OPCS runs in $\mathcal{O}(n^{10})$ time. We further optimize the running time for orthogonal polygons with $n$ vertices and a small number of knobs $k$, by designing a recursive algorithm with running time $\mathcal{O}(n^{2}+k^{10}\cdot n)$. This gives us an $\mathcal{O}(n^{2})$ algorithm for solving OPCS on orthogonally convex polygons. We also provide a correct proof for the NP-hardness of OPCSH; this is a novel result as the proof claimed in the works of Aupperle et. al [4] incorrectly reduce from a polynomial-time solvable problem. A natural future direction is to study the problem with respect to covering of orthogonal polygons with other geometric objects like triangles, line segments, orthogonally convex polygons etc. and try to obtain an algorithm whose running time is polynomial in the number of vertices of the input polygon and not on the total number of interior lattice points of the input polygon. Another interesting question directly related to our current work would be to find tight bounds for the number of rec-packs required to produce a minimum covering of an orthogonal polygon; our paper only proves an upper bound of $\mathcal{O}(n^{2})$ rec-packs required to produce a minimum covering of an entire polygon.